# The DSA Transparency Database: Auditing Self-reported Moderation Actions by Social Media

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; Sometime 01–02, 2024; Neverlandisbn: 978-1-4503-XXXX-X/YY/ZZccs: Human-centered computing Empirical studies in collaborative and social computingccs: Information systems Social networking sitesccs: Social and professional topics Governmental regulations
@@ 1. Introduction
Content moderation refers to the strategies and processes through which an online platform ensures that the content and behaviors it hosts adhere to the platform’s self-defined standards and guidelines, and to legal regulations (Gillespie, 2018). Moderation interventions are the specific actions through which platforms enforce such strategies and processes. Currently, a wide array of interventions are adopted, which involve removing content and users (Jhaver et al., 2021), demoting, restricting, or otherwise reducing their visibility (Le Merrer et al., 2021; Trujillo and Cresci, 2022), showing users warning messages (Katsaros et al., 2022) or attaching warning labels to problematic content (Zannettou, 2021), and even acting on monetary rewards (Ma and Kou, 2022).
However, despite their importance in maintaining safe and healthy online environments, the application of these interventions often sparks concerns. The controversies are multifaceted, such as considering the potential misuse of removals and reduced visibility decisions as a form of censorship that threatens users’ freedom of speech (Myers West, 2018). Furthermore, transparency in content moderation is often questioned, as the processes behind moderation decisions —whether executed by black-box AI algorithms or human moderators— are sometimes perceived as opaque, or lacking clarity and consistency (Gillespie, 2020). Issues of explainability, fairness, and the susceptibility of automated systems to biases further contribute to the ongoing debate (Nogara et al., 2023), highlighting the delicate balance that platforms must strike between upholding community standards and legal obligations, and respecting user rights in the digital sphere (Grimmelmann, 2015).
@@ 1.1. The Digital Services Act
  • Article 33 designates those platforms with more than 45M users in the EU as either very large online platforms (VLOPs) or very large online search engines (VLOSEs). Due to their strategic role, these must comply with the most stringent rules of the DSA. On April 2023, the EU designated an initial set of 17 VLOPs and 2 VLOSEs, [^1] https://ec.europa.eu/commission/presscorner/detail/en/IP_23_2413 whereas on December 2023 it identified a second set of 3 VLOPs focused on adult content. [^2] https://ec.europa.eu/commission/presscorner/detail/en/ip_23_6763
@@
  • Article 17 obliges VLOPs and VLOSEs to submit a clear and specific statement of reasons (SoR) for each of their moderation interventions. Each SoR must provide detailed and timely information on the intervention, its legal ground, and the content to which it was applied.
@@ 1.2. Research Questions and Contributions
Herein, we carry out an in depth quantitative analysis of very large social media moderation practices, as reflected by their SoR submissions to the DSA-TDB. We analyze all 353.12M SoRs submitted by eight platforms during the first 100 days of the DSA-TDB, providing an early yet thorough look into (i) the overall frequency of moderation interventions by very large social media, (ii) their grounds for decision, (iii) the types of applied restrictions, (iv) the types of moderated content, (v) the timeliness of their interventions, and (vi) their use of automation in content moderation. Figure 1 depicts this 100-day period of study (orange-colored) in the context of relevant events related to the DSA-TDB. In carrying out these analyses, we specifically focus on answering to the following research questions.
@@ 2.2. Platform transparency on content moderation
Social media platforms, crucial as they are in shaping public discourse, face scrutiny over their content moderation practices. Balancing freedom of expression with digital civil discourse is challenging (DeNardis and Hackl, 2015; Gillespie, 2018), and criticism arises due to a perceived lack of transparency and biases in moderation (Jhaver et al., 2019; Myers West, 2018). Far-right users claim censorship (Jahanbakhsh et al., 2021; Myers West, 2018), while left-leaning users point to policy enforcement inconsistencies (Shen and Rosé, 2019). In addition, marginalized communities often face disproportionate moderation (Haimson et al., 2021; Thach et al., 2022). The evolution from manual to automated moderation (Zhang et al., 2021; Annamoradnejad and Habibi, 2023) further complicates matters, with mixed user opinions on their fairness and bias (Lyons et al., 2022; Gonçalves et al., 2023; Nogara et al., 2023). Therefore, transparent content moderation practices by online platforms are of paramount importance. In this regard, the DSA provides three important mechanisms: periodic transparency reports, granting researchers access to platform data, and the new centralized transparency database —the subject of the present study.
@@
Having access to data posted on social media through specific APIs or historical repositories is also important toward moderation transparency. Arguably, the then free Twitter Researcher API (Chen et al., 2022) and Reddit Pushshift archive (Baumgartner et al., 2020) were among the most open and widely used data for social media analysis. For several years, these enabled researchers to investigate the effects of moderation policies of their respective platforms. For instance, several studies have investigated the impact of major moderation interventions on user behavior in Reddit, at the community (Trujillo and Cresci, 2022) and individual levels (Trujillo and Cresci, 2023), inside the platform as well as across platforms (Horta Ribeiro et al., 2021). Interestingly, in several cases —an particularly in the case of political communities— such moderation interventions had mostly mixed effects (Chandrasekharan et al., 2022; Shen and Rosé, 2022).
@@
In early 2023, due to changes in management and in part to the unauthorized use of social media content to train generative AI models, the Twitter API ceased to be free [^4] https://twitter.com/XDevelopers/status/1621026986784337922 (even for research) and Reddit severely restricted access to its API. [^5] https://redd.it/12qwagm/ Indeed, both platforms, as well as others, changed their terms of service with respect to data access and collection. Consequently, many social media analysis projects were halted, significantly hindering the transparency and accountability of social media. In this respect, the DSA includes a provision for granting vetted researchers access to VLOPs and VLOSEs data, provided that certain conditions are met (European Parliament and Council, 2022). The overarching goal is that of understanding how society is shaped by these platforms, as well as supporting their oversight of DSA compliance. Some of the very large platforms have just recently set up these new mechanisms to access EU content, hence the effect of such provision on online moderation research remains to be seen.
@@ 2.3. Precursory analyses of the DSA Transparency Database
The DSA-TDB, an unprecedented potential data trove regarding moderation actions, has piqued much scholarly interest since its announcement. Even at this early stage, two non-peer-reviewed works provided preliminary analyses of the initial data submitted by large social media platforms to the database. The study by Dergacheva et al. (2023) only took into consideration submissions made during the very first day of the database. However, in spite of the extremely limited breadth of the analysis, they found much heterogeneity in the reported degrees of automation, applied visibility measures, and types of content moderated by various online platforms on a single day. In a work more similar to our present contribution, Drolsbach and Pröllochs (2023) examined the first 60 days of data from the database, revealing relevant disparities in moderation practices across different platforms. Their findings underscored the predominance of automated moderation, with X (formerly Twitter) being the main exception, and highlighted some inconsistencies in implementing the DSA guidelines. Differently from our study, the work by Drolsbach and Pröllochs (2023) is mostly descriptive and it does not systematically take into consideration the differences among platforms. For instance, a regression analysis was conducted regarding the use of automated means for moderation, which however conflated different levels of automation and data from various platforms. Given that TikTok represented more than 60% of the data, their resulting model is heavily influenced by the self-reported practices of a single platform, with very different dynamics compared to the other seven social media.
@@
On the other hand, our research presents a different approach to the database and includes several aspects not covered in any of the previous analyses. Firstly, we investigate the initial 100 days of the database, providing a much more comprehensive and robust time frame. Secondly, we conduct a thorough platform-wise analysis of moderation actions, taking into consideration the potential distinctiveness of each platform. Thirdly, our study provides a deeper assessment of the data submitted to the database, specifically addressing questions about the adequacy, consistency, and timeliness of submitted SoRs. Lastly, we also carry out a systematic coherence cross-check of the database against each platforms’ public DSA Transparency Reports, which allows to identify discrepancies and inaccuracies in the reported data. Our approach not only uncovers worrisome inconsistencies, but also critically evaluates the effectiveness and limitations of the current database schema, while suggesting possible enhancements.
@@ 3. Overview of the DSA Transparency Database
Out of the 17 VLOPs initially designated by the DSA, 16 have submitted at least one SoR. Wikipedia is the only VLOP that did not make any submission to the DSA-TDB during the initial 100-day period. In this time, a total of 1.4861B SoRs were submitted. The largest submitter was by far Google Shopping, with more than a billion SoRs (67.8% of the total), and the smallest was Zalando, with merely 9 SoRs submitted. Given our focus on social media —platforms centered on user generated content and social networking— we selected the following eight platforms of interest for this study: Facebook, Instagram, LinkedIn, Pinterest, Snapchat, TikTok, X (formerly Twitter), and YouTube. With 353.12M SoRs, these platforms account for 23.8% of the whole database.
@@ 4. Analyses and Results
Upon exploration of the database, we realized that platforms had varying approaches in how they filled in and mapped the attributes of their SoR submissions, especially in terms of mandatory and optional fields. For instance, only 3 out of 16 VLOPs indicate SoR content language —an optional attribute— with none of the social media doing so. Therefore, in this section we present only the aspects which we considered more pertinent in answering our research questions, namely: grounds for the decision on an infringement, types of restriction decisions, type of moderated content, timeliness of intervention application and communication to the DSA-TDB, and the use of automated means for moderation. As a final validation step, we compare the data submitted by each social media to the DSA-TDB with the respective claims from their DSA Transparency Reports, so as to detect possible inconsistencies.
@@ 4.1. Grounds for Decision
There are several attributes to indicate the legal or contractual grounds relied on in taking a moderation intervention decision, and to codify the infringement type. We first considered two grounds attributes: a mandatory decision ground (illegal, incompatible), and an optional attribute to indicate that a content is both incompatible with the terms of service and illegal. We then analyzed three attributes to codify infringements: category (mandatory high-level attribute from a predefined set of 14 categories), category addition (optional list from the same set of categories), and category specification (optional keyword list from a set of 55 predefined values).
Concerning the decisions grounds, for almost every platform incompatible content was the overwhelming majority (more than 99%), up to all instances in the case of LinkedIn. The exception was X, which indicated illegal content for all of its SoRs. Curiously, only Snapchat indicated an overlap of incompatible and illegal content in 12.2% of their SoRs. None of the other social media platforms used this attribute.
@@
    those in which scope of platform service [^7] This is a generic category to indicate content moderated because deemed outside of the scope of the platform service. was the most common infringement (YouTube, Facebook, LinkedIn, TikTok, and Instagram);
@@
4.2. Types of Restriction Decisions
There are four main types of restriction decisions: account (suspension, termination), monetary (suspension, termination, other), provision (partial/total suspension/termination of the service), and visibility (restriction for the content). Platforms must indicate a single sub-type of one of the four types, but these types are not mutually exclusive. In other words, a SoR can contain multiple types of restrictions, with at least one of them being mandatory. Nonetheless, all of the social media platforms only set a single restriction type. As detailed in Table 2, most were in visibility (79.3% across platforms), followed by account (17.6%) and provision (3.1%). Surprisingly, there was not a single SoR with a monetary restriction decision. For visibility sub-types, shown in Figure 3, we noticed a similar platform grouping to that of infringement categories:

### v4

@@ Abstract.
copyright: rightsretainedjournal: PACMHCIjournalyear: 2025journalvolume: 9journalnumber: 2article: CSCW187publicationmonth: 4doi: 10.1145/3711085ccs: Human-centered computing Empirical studies in collaborative and social computingccs: Information systems Social networking sitesccs: Social and professional topics Governmental regulations
@@ 1. Introduction
Content moderation refers to the strategies and processes through which an online platform ensures that the content and behaviors it hosts adhere to the platform’s self-defined standards and guidelines, and to legal regulations (Gillespie, 2018). Moderation interventions are the specific actions through which platforms enforce such strategies and processes. Currently, a wide array of interventions are adopted, which involve removing content and users (Jhaver et al., 2021; Tessa et al., 2024), demoting, restricting, or otherwise reducing their visibility (Le Merrer et al., 2021; Trujillo and Cresci, 2022), showing users warning messages (Katsaros et al., 2022) or attaching warning labels to problematic content (Zannettou, 2021), and even acting on monetary rewards (Ma and Kou, 2022).
However, despite their importance in maintaining safe and healthy online environments, the application of these interventions often sparks concerns. The controversies are multifaceted, such as considering the potential misuse of removals and reduced visibility decisions as a form of censorship that threatens users’ freedom of speech (Myers West, 2018). Furthermore, transparency in content moderation is often questioned, as the processes behind moderation decisions —whether executed by black-box AI algorithms or human moderators— are sometimes perceived as opaque, or lacking clarity and consistency (Gillespie, 2020). Issues of explainability, fairness, and the susceptibility of automated systems to biases further contribute to the ongoing debate (Nogara et al., 2024), highlighting the delicate balance that platforms must strike between upholding community standards and legal obligations, and respecting user rights in the digital sphere (Grimmelmann, 2015).
@@ 1.1. The Digital Services Act
  • Article 33 designates those platforms with more than 45M users in the EU as either very large online platforms (VLOPs) or very large online search engines (VLOSEs). Due to their strategic role, these must comply with the most stringent rules of the DSA. As of July 2024, the EU has designated 22 VLOPs and 2 VLOSEs. [^1] https://digital-strategy.ec.europa.eu/en/policies/list-designated-vlops-and-vloses
@@
  • Article 17 obliges VLOPs and VLOSEs to submit a clear and specific statement of reasons (SoR) for each of their moderation interventions. Each SoR must provide detailed and timely information on the intervention, its legal ground, and the content to which it was applied. Since February 2024 this article applies to all online platforms independently of their size.
@@
[…]
17 February 2023 was the deadline for platforms to publish user numbers. On 25 April the European Commission designated 17 VLOPs and 2 VLOSEs. On 25 September 2023 the DSA Transparency Database begins operating. 5 November was the deadline to publish transparency reports for the first semester of 2023. On 21 December the European Commision designated another 3 VLOPs, all of which focus on adult content. From 17 February 2024 the DSA rules are applied to all regulated online platforms, not only those deemed very large.
@@ 1.2. Research Questions and Contributions
Herein, we carry out an in depth quantitative analysis of very large social media moderation practices, as reflected by their SoR submissions to the DSA-TDB. We analyze all 353.12M SoRs submitted by eight platforms during the first 100 days of the DSA-TDB, providing an early yet thorough look into (i) the overall frequency of moderation interventions by very large social media, (ii) their grounds for decision, (iii) the types of applied restrictions, (iv) the types of moderated content, (v) the timeliness of their interventions, and (vi) their use of automation in content moderation. Figure 1 depicts this 100-day period of study (orange-colored) in the context of relevant events related to the DSA-TDB. In carrying out these analyses, we specifically focus on answering the following research questions.
@@ 2.2. Platform Transparency on Content Moderation
Social media platforms, crucial as they are in shaping public discourse, face scrutiny over their content moderation practices. Balancing freedom of expression with digital civil discourse is challenging (DeNardis and Hackl, 2015; Gillespie, 2018), and criticism arises due to a perceived lack of transparency and biases in moderation (Jhaver et al., 2019; Myers West, 2018). Far-right users claim censorship (Jahanbakhsh et al., 2021; Myers West, 2018), while left-leaning users point to policy enforcement inconsistencies (Shen and Rosé, 2019). In addition, marginalized communities often face disproportionate moderation (Haimson et al., 2021; Thach et al., 2022). The evolution from manual to automated moderation (Zhang et al., 2021; Annamoradnejad and Habibi, 2023) further complicates matters, with mixed user opinions on their fairness and bias (Lyons et al., 2022; Gonçalves et al., 2023; Nogara et al., 2024). Therefore, transparent content moderation practices by online platforms are of paramount importance. In this regard, the DSA provides three important mechanisms: periodic transparency reports, granting researchers access to platform data, and the new centralized transparency database —the subject of the present study.
@@
Having access to data posted on social media through specific APIs or historical repositories is also important toward moderation transparency. Arguably, the then free Twitter Researcher API (Chen et al., 2022) and Reddit Pushshift archive (Baumgartner et al., 2020) were among the most open and widely used data for social media analysis. For several years, these enabled researchers to investigate the effects of moderation policies of their respective platforms. For instance, several studies have investigated the impact of major moderation interventions on user behavior in Reddit, at the community (Trujillo and Cresci, 2022) and individual levels (Trujillo and Cresci, 2023), inside the platform as well as across platforms (Horta Ribeiro et al., 2021). Interestingly, in several cases —and particularly in the case of political communities— such moderation interventions had mostly mixed effects (Chandrasekharan et al., 2022; Shen and Rosé, 2022).
@@
In early 2023, due to changes in management and in part to the unauthorized use of social media content to train generative AI models, the Twitter API ceased to be free [^3] https://twitter.com/XDevelopers/status/1621026986784337922 (even for research) and Reddit severely restricted access to its API. [^4] https://redd.it/12qwagm/ Indeed, both platforms, as well as others, changed their terms of service with respect to data access and collection. Consequently, many social media analysis projects were halted, significantly hindering the transparency and accountability of social media. In this respect, Article 40 of the DSA establishes a provision for granting vetted researchers access to VLOPs and VLOSEs data, provided that certain conditions are met (European Parliament and Council, 2022). The overarching goal is that of understanding how society is shaped by these platforms, as well as supporting their oversight of DSA compliance. However, some of the very large platforms have just recently set up the mechanisms to access EU content, while independent investigations documented widespread hurdles at accessing platform data under Article 40 of the DSA, raising serious concerns on platform compliance with the regulation (Jaursch et al., 2024). Therefore, the effect of such provision on online moderation research remains to be seen.
@@ 2.3. Precursory Analyses of the DSA Transparency Database
The DSA-TDB, an unprecedented potential data trove regarding moderation actions, has piqued much scholarly interest since its announcement. Even at this early stage, a few works have provided preliminary analyses of the initial data submitted by large social media platforms to the database. Kaushal et al. (2024) performed a multidisciplinary legal and computational analysis of a data sample of 131M SoRs from all VLOPs in the database, including shopping platforms, social media, app stores, and service platforms. They concluded that while the database provides some insight into platform moderation actions, its reliance on self-reported data limits transparency and legal certainty, revealing inconsistencies, potential manipulation, and insufficient compliance with the DSA’s objectives, necessitating complementary transparency mechanisms and robust enforcement by the European Commission. A different study investigated the degree of variability of content moderation practices for the main social media platforms across the EU member states, with a particular focus on the use of automated moderation (Papaevangelou and Votta, 2024). By analyzing four months of data, the study found inconsistent information and a great deal of heterogeneity between the different platforms, and even for certain platforms across different EU member states. The study by Dergacheva et al. (2023) only took into consideration submissions made during the very first day of the database. However, in spite of the extremely limited breadth of the analysis, they found much heterogeneity in the reported degrees of automation, applied visibility measures, and types of content moderated by various online platforms on a single day.
In a work more similar to our present contribution, Drolsbach and Pröllochs (2023) examined the first 60 days of data from the database, revealing relevant disparities in moderation practices across different platforms. Their findings underscored the predominance of automated moderation, with X (formerly Twitter) being the main exception, and highlighted some inconsistencies in implementing the DSA guidelines. Differently from our study, the work by Drolsbach and Pröllochs (2023) is mostly descriptive and it does not systematically take into consideration the differences among platforms. For instance, a regression analysis was conducted regarding the use of automated means for moderation, which however conflated different levels of automation and data from various platforms. Given that TikTok represented more than 60% of the data, their resulting model is heavily influenced by the self-reported practices of a single platform, with very different dynamics compared to the other seven social media.
@@
On the other hand, our research presents a different approach to the database and includes several aspects not covered in any of the previous analyses. Firstly, we investigate the initial 100 days of the database, providing a much more comprehensive and robust time frame. Secondly, we conduct a thorough platform-wise analysis of moderation actions, taking into consideration the potential distinctiveness of each social media platform. Thirdly, our study provides a deeper assessment of the data submitted to the database, specifically addressing questions about the adequacy, consistency, and timeliness of submitted SoRs. Lastly, we also carry out a systematic coherence cross-check of the database against each platforms’ public DSA Transparency Reports, which allows to identify discrepancies and inaccuracies in the reported data. Our approach not only uncovers worrisome inconsistencies, but also critically evaluates the effectiveness and limitations of the current database schema, while suggesting possible enhancements.
@@ 3. Overview of the DSA Transparency Database
Out of the 17 VLOPs initially designated by the DSA, [^6] https://ec.europa.eu/commission/presscorner/detail/en/IP_23_2413 16 have submitted at least one SoR. Wikipedia is the only VLOP that did not make any submission to the DSA-TDB during the initial 100-day period. In this time, a total of 1.4861B SoRs were submitted. The largest submitter was by far Google Shopping, with more than a billion SoRs (67.8% of the total), and the smallest was Zalando, with merely 9 SoRs submitted. Given our focus on social media —platforms centered on user generated content and social networking— we selected the following eight platforms of interest for this study: Facebook, Instagram, LinkedIn, Pinterest, Snapchat, TikTok, X (formerly Twitter), and YouTube. With 353.12M SoRs, these platforms account for 23.8% of the whole database.
@@ 4. Analyses and Results
Upon exploration of the database, we realized that platforms had varying approaches in how they filled in and mapped the attributes of their SoR submissions, especially in terms of mandatory and optional fields. For instance, only 3 out of 16 VLOPs indicate SoR content language —an optional attribute— with none of the social media doing so. Therefore, in this section we present only the aspects that we considered more pertinent in answering our research questions, namely: grounds for the decision on an infringement, types of restriction decisions, type of moderated content, timeliness of intervention application and communication to the DSA-TDB, and the use of automated means for moderation. When presenting and discussing these aspects, we make explicit reference to the corresponding section of the official documentation [^7] https://transparency.dsa.ec.europa.eu/page/documentation at the time of writing. This allows us to map each analysis to the corresponding field(s) in the dataset that we utilized, providing a clear reference to the documentation to enhance understanding of the field(s) semantics. Furthermore, by comparing our findings with the documentation, we gain valuable context on how the platforms have used the DSA-TDB fields, in relation to their intended or expected use. As a final validation step, we compare the data submitted by each social media to the DSA-TDB with the respective claims from their DSA Transparency Reports, so as to detect possible inconsistencies.
@@ 4.1. Grounds for Decision
There are several attributes to indicate the legal or contractual grounds relied on in taking a moderation intervention decision, and to codify the infringement type. We first considered two grounds attributes (§11 Decision Grounds in the official documentation): a mandatory decision ground (illegal, incompatible), and an optional attribute to indicate that a content is both incompatible with the terms of service and illegal. We then analyzed three attributes to codify infringements (§16 Category & Specification): category (mandatory high-level attribute from a predefined set of 14 categories), category addition (optional list from the same set of categories), and category specification (optional keyword list from a set of 55 predefined values). Concerning the decisions grounds, for almost every platform incompatible content was the overwhelming majority (more than 99%), up to all instances in the case of LinkedIn. The exception was X, which indicated illegal content for all of its SoRs. Curiously, only Snapchat indicated an overlap of incompatible and illegal content in 12.2% of their SoRs. None of the other social media platforms used this attribute.
[…]
A heatmap shows that platforms can be grouped into two groups with similar distributions, plus a single-platform group that does not share a similar distribution to others in terms of infringement categories. A first group is composed on Pinterest and X, in which sexual and violent content are the most common infringement categories. A second group is composed of YouTube, Facebook, LinkedIn, TikTok, and Instagram, for which the category scope of platform service was the most common. A third group is composed of Snapchat, which unlike the other platforms, does not present a dominant category.
@@
    those in which scope of platform service [^8] This category indicates content moderated because deemed outside of the scope of the platform service. It includes content with age-specific restrictions, with geographical or language requirements, disallowed goods and services, and nudity. was the most common infringement (YouTube, Facebook, LinkedIn, TikTok, and Instagram);
@@ 4.2. Types of Restriction Decisions
There are four main types of restriction decisions (§3 The type of restriction(s) imposed): account (suspension, termination), monetary (suspension, termination, other), provision (partial/total suspension/termination of the service), and visibility (restriction for the content). Platforms must indicate a single sub-type of one of the four types, but these types are not mutually exclusive. In other words, a SoR can contain multiple types of restrictions, with at least one of them being mandatory. Nonetheless, all of the social media platforms only set a single restriction type. As detailed in Table 2, most were in visibility (79.3% across platforms), followed by account (17.6%) and provision (3.1%). Surprisingly, there was not a single SoR with a monetary restriction decision. For visibility sub-types, shown in Figure 3, we noticed a similar platform grouping to that of infringement categories:
@@
[…]
A heatmap shows that platforms can be grouped into two groups with similar distributions, plus a single-platform group that does not share a similar distribution to others in terms of decisions on visibility. A first group is composed on Pinterest and X, in which other is the most common. A second group is composed of TikTok, LinkedIn, Instagram, Facebook, and YouTube, for which the visibility content removed was the most common. A third group is composed of Snapchat, which has content disabled as the most common visibility.